1. Teal Companys unadjusted trial balance at December 31, 2020, included the following accounts. Credit Accounts receivable5. The following accounts were taken from Sarasota Inc.s trial balance at December 31, 2020. Debit Credit $687,000 Net credi
Answer #1

Requirement 1 Bad debt expense for 2020 = (8%x49,900)+7,790 =11,782 Requirement 2 Net realizable value = 1,791,100-182,680 =1